# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v1

@@ 6.2 Performance
Table 3: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods. All methods employ 50% density in both prefill and decode stages, except for H2O, which uses dense prefill.
Model Attention 4k 8k 16k 16k 0-4k 4-8k 8-16k 8-12k
Vicuna-7B Original 1.00 0.98 0.62 5.93 37.91 33.82 32.54 3.79
H2O 0.86 0.68 0.35 5.42 36.23 32.74 31.81 3.94
StreamingLLM 0.43 0.16 0.08 4.72 30.53 33.28 31.70 4.48
MoA 1.00 0.97 0.57 5.61 37.04 32.90 31.94 3.75
Vicuna-13B Original 0.99 0.98 0.44 5.83 42.25 39.52 35.93 3.62
H2O 0.88 0.76 0.28 5.66 41.63 38.02 34.75 3.80
StreamingLLM 0.65 0.49 0.33 5.43 30.65 33.07 32.68 4.10
MoA 0.99 0.93 0.49 7.16 41.73 38.88 35.69 3.62
Llama3-8B Original 0.99 0.99 0.97 17.49 44.27 43.53 43.26 4.52
H2O 0.94 0.89 0.88 16.03 43.46 43.01 42.50 4.63
StreamingLLM 0.68 0.55 0.52 11.16 37.20 38.02 39.43 4.79
MoA 0.99 1.00 1.00 17.46 43.07 42.75 43.09 4.49
Overall Performance. As detailed in Table 3, MoA outperforms state-of-the-art sparse attention methods and achieves comparable performance to the original dense model at 50% density. We calculate the relative performance drop of the compressed model over the original one. For retrieval tasks, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87% and 44% for StreamingLLM and H2O. On average, the relative accuracy drop for MoA is under 1%, while it is much higher for StreamingLLM and H2O at 51% and 20%, respectively. In benchmarks using LV-Eval and LongBench, MoA shows only a 5% and 3% maximum relative drop in scores, respectively, compared to steep declines for StreamingLLM at 36% and 27%. H2O exhibits maximum drops of 9% and 4% with higher efficiency costs. Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while StreamingLLM and H2O exhibit 13% and 4% increases. This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3.
Figure 4: Retrieval accuracy tests on LongEval. With Vicuna-7B: (a) varies input lengths and densities to show effective context lengths across attention spans, (b) sets input length at 8k and shows retrieval accuracy across different densities. With Llama3-8B: (c) sets density at 50% and shows retrieval accuracy across increasing input lengths.
Long-Context Retrieval. We investigate the ultimate effective context length with retrieval tests on LongEval [35]. As shown in Figure 4(a), MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model. Detailed data for effective context length calculation are in Appendix B.1.2. Figure 4(b) shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively. Figure 4(c) further investigates MoA’s ability to expand and generalize to extreme input lengths. By profiling within 8k lengths, MoA extends its capability to 60k inputs while maintaining over 90% retrieval accuracy at 50% density.
Figure 5: (a) LV-Eval and (b) LongBench scores for different attention methods at 50% density, tested on Vicuna-7B and 13B models. Scores normalized against the original dense model.
Long-Context Understanding. Figure 5 illustrates the performance comparison across long-context benchmarks LV-Eval and LongBench. MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost. In contrast, StreamingLLM displays inconsistent performance: although it occasionally surpasses the original model, it generally suffers noticeable degradation.
6.3 Efficiency
Table 4: Efficiency analysis of different frameworks on 7B and 13B models. H2O and MoA use 50% density. GPU memory evaluated with batch sizes 8 (7B model) and 4 (13B model). Decode throughput evaluated at the maximum batch capacity of an A100-80GB GPU.
Memory (GB) Throughput (tok/s)
Size Framework 4k 8k 16k 4k 8k 16k
7B FlashAttn2 28.5 44.4 76.3 134.6 66.9 32.9
H2O 36.9 OOM OOM 754.9 296.3 51.7
MoA 22.7 32.9 53.5 897.7 436.7 206.4
13B FlashAttn2 36.8 49.2 74.0 81.3 40.8 19.8
H2O 40.4 77.9 OOM 330.2 138.2 37.4
MoA 32.0 39.6 55.0 471.5 222.6 108.3
@@ 6.3 Efficiency
We use Huggingface [66] with FlashAttention [13] as the baseline and implement MoA based on it. As shown in Table 5, with an average 50% density, MoA reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 × across various model sizes and input lengths. Moreover, MoA boosts decode throughput by $\times5$ to 6.7×6.7$\times$6.7 ×, primarily attributed to a static-size KV-Cache (3.0$\approx$ 3.0$\times$ 3.0 ×), reduced attention computations (1.5$\approx$ 1.5$\times$ 1.5 ×) and increased batch sizes (1.4$\approx$ 1.4$\times$ 1.4 ×). Compared with H2O that uses dense prefill and requires additional cost for dynamic mask generation, MoA achieves $1.2$ to $4.0$ throughput at the same 50% density. Compared to the highly optimized vLLM framework [33], MoA still achieves a 1.4×1.4$\times$1.4 × to 1.5×1.5$\times$1.5 × increase in throughput. More efficiency results are shown in Appendix B.2.
@@ 6.4 Rules Discovered by MoA
We investigate MoA’s elastic rules for each head. As shown in Figure 8, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions [61] and layer sensitivities [71]. Conversely, in the final layers, most heads require low density, while few need high density. Figure 9 shows that layers with lower average mask density typically display a wider range of densities among heads, confirming the need for heterogeneity within the same layer. Further details and quantified insights are presented in Appendix C.
@@ 7 Conclusion and Future Work
MoA automates the selection of heterogeneous elastic masks for each attention head and input length, significantly extending the effective context length of LLMs by 3.9×3.9$\times$3.9 ×. It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × and increases throughput to over 5×5$\times$5 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
Limitations and Future Work. Under an extremely low-density budget, MoA fails to maintain good performance. Designing a dynamic MoA method has the potential to address this issue, which we leave for future work. To further enhance MoA’s efficiency, system, and kernel-level optimizations such as kernel fusion [27] and KV-Cache management [33] could be integrated. Using non-linear elastic rules with bounded attention spans is also worth exploring. Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization [40, 68, 37].
@@ A.1 Main Setup
Baselines. In the setup for our experiment, we adhere to specific configurations outlined in the respective papers. In the case of StreamingLLM [69], the initial four tokens remain unmasked, serving as the attention sink. For H2O [75], we ensure the same number of heavy hitter tokens and recent tokens. Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction. StreamingLLM and MoA use sparse prefill.
@@
Models and Benchmarks. Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k [8] can only take in 16k context length, we use the 16k split of LV-Eval benchmark [70], truncating the input to 15500 for model input. For the LongBench benchmark [4], we use the LongBench-E split, which features a balanced number of data items at every length level. The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens. We test each split using the input length truncation thresholds of 3,500, 7,500, and 15,500 tokens, respectively.
@@
Perplexity Evaluation. We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper [15], MultiNew  [17], TREC [38, 28] and LCC [45], representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM. Following LongBench, the data items are organized as question-answer pairs. The questions and answers are written by humans and come with the dataset. The perplexity is calculated solely on the answer part of the data, demonstrating the model’s coherence in responding to user requests.
@@
Validation Dataset. The validation dataset is used to select the optimal compression plan among the Pareto front solutions during the optimization step. The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets. $50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper [15], MultiNew  [17], TREC [38, 28] and LCC [45]. The additional $50$ data items from the LongEval [35] dataset are also added to validate the retrieval ability. For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
@@
MoA Settings. MoA uses the block sparse attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block. The first block of tokens is not masked as the attention sink. For the profile stage, we use the MultiNews [17] calibration dataset with model response as supervision, as described in Section 5. We use $50\times 3$ data items at 2k, 4k, 8k lengths. The data items are padded to their corresponding length level in order to ensure a unified shape of attention influence tensors for each length level. We adopt block granularity during the profiling stage, calculating the average attention influence within each block to represent the block’s overall influence. The optimization is done with the multi-objective optimization at the same set of lengths. Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
@@ A.2 Efficiency Experiment Setup
For MoA, The implementation is based on Huggingface Transformers. During the prefill stage, we use the sparse triton [58] kernel designed by us with block size 64. During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules. Thanks to our fixed sliding-window span during the decode stage, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache. Then, the dense FlashAttention [13] decoding is applied with the updated KV-Cache.
For H2O, we use its official efficient implementation, which is based on Flexgen [54]. Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated. It makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention. Therefore, H2O is easy to get OOM (Out-Of-Memory) with large prefill length and increased batch size.
@@
Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible. However, the official efficient implementation of H2O based on Flexgen only supports OPT [74]. Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 5 for comparison.

### v2

@@ 6.3 Performance
Figure 4: Accuracy-throughput trade-offs of various attention methods at different densities, tested on Vicuna-7B with 8k input length using one A100-80GB GPU on the LongEval dataset.
@@
Table 4: Comparative analysis of retrieval accuracy, LV-Eval scores, LongBench scores, and perplexity for various models with different attention methods. All sparse methods employ 50% density in decode stage. H2O uses dense prefill, while StreamingLLM, InfLLM and MoA use sparse prefill. InfLLM for 70B model is excluded due to OOT issues.
Model Attention 4k 8k 16k 16k 0-16k 8-12k
Vicuna-7B Original 1.00 0.98 0.62 5.93 34.76 3.79
H2O 0.86 0.68 0.35 5.42 33.59 3.94
InfLLM 0.67 0.57 0.26 5.13 32.97 4.07
StreamingLLM 0.43 0.16 0.08 4.72 31.84 4.48
MoA 1.00 0.97 0.57 5.61 33.96 3.75
Vicuna-13B Original 0.99 0.98 0.44 5.83 39.23 3.62
H2O 0.88 0.76 0.28 5.66 38.13 3.80
InfLLM 0.70 0.53 0.27 6.80 37.13 4.07
StreamingLLM 0.65 0.49 0.33 5.43 32.13 4.10
MoA 0.99 0.93 0.49 7.16 38.77 3.62
Llama3-8B Original 0.99 0.99 0.97 17.49 43.69 4.52
H2O 0.94 0.89 0.88 16.03 42.99 4.63
InfLLM 0.65 0.59 0.37 14.44 42.43 4.68
StreamingLLM 0.68 0.55 0.52 11.16 38.22 4.79
MoA 0.99 1.00 1.00 17.46 42.97 4.49
Llama3-70B Original 1.00 0.99 0.93 24.51 49.10 3.67
H2O 0.93 0.91 OOM OOM OOM OOM
StreamingLLM 0.20 0.15 0.04 17.45 42.53 4.26
MoA 1.00 1.00 0.94 23.65 47.79 3.75
MoA outperforms state-of-the-art sparse attention methods across various model sizes and benchmarks, achieving comparable performance to the original dense model at 50% density.
Long-Context Retrieval. As shown in Table 4, MoA demonstrates a maximum of 8% relative accuracy drop (calculated as $\max\{1-\text{Acc.}_{\text{MoA}}/\text{Acc.}_{\text{Original}}\}$ across three lengths and LLMs), significantly less than the 87%, 58% and 44% for StreamingLLM, InfLLM and H2O. On average, the relative accuracy drop for MoA is under 1%, much less than others at 51%, 41% and 20%, respectively. Figure 5(a) shows that MoA retains over 90% retrieval accuracy up to 60k lengths, equaling the dense model’s effective context length. Note that it is done within 8k profiling and 12k validation. In contrast, the effective context lengths for H2O, InfLLM, and StreamingLLM are only 8k, <4k, and <4k, respectively. Appendix B.1.2 shows that MoA extends its effective context to approximately 3.9×3.9$\times$3.9 × the average KV-Cache length.
Long-Context Understanding. As shown in Table 4, MoA minimizes the maximum relative performance drop in LV-Eval and LongBench benchmarks to only 5% and 3%, respectively—much lower than the 36% and 18% experienced by StreamingLLM. H2O and InfLLM show maximum relative drops of 9%-17% and 3%-5% with higher efficiency costs. Similar trends show in perplexity tests, where MoA maintains less than 1% relative perplexity increase, while others exhibit 4%-13% increases. This trend holds for other densities, as shown in Appendices B.1.1 and B.1.3. Figure 9 and Table 8 further details the score with different tasks. MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost. In contrast, StreamingLLM and InfLLM display inconsistent performance: it sometimes surpasses the original model in some tasks, while suffering noticeable degradation in others.
Longer-Context Generalization. By compressing within 12k lengths, MoA effectively generalizes to lengths of 32k-256k, as shown in Figure 5(b). At the extended lengths, MoA outperforms both InfLLM and StreamingLLM by 1.93.3×1.9-3.3$\times$1.9 - 3.3 × in retrieval accuracy and 1.21.4×1.2-1.4$\times$1.2 - 1.4 × in LV-Eval scores, demonstrating comparable performance to the original dense model.
(a) Retrieval accuracy and the effective context length (arrow).
Retrieve Acc. $\uparrow$ LV-Eval $\uparrow$
Attention 32k 64k 128k 256k 32k 64k 128k
Original 0.98 0.93 0.76 0.37 16.74 15.39 14.71
InfLLM 0.43 0.32 0.25 OOT 14.22 12.17 OOT
StreamingLLM 0.52 0.48 0.41 0.25 12.38 11.45 11.94
MoA 1.00 0.92 0.83 0.46 17.07 15.13 14.14
(b) Retrieval accuracy and LV-Eval score at longer lengths
Figure 5: Comparative analysis at extended sequence lengths with different attention methods using Llama3-8B model. All methods employ 50% density in both prefill and decode stages.
6.4 Efficiency
Table 5: Runtime efficiency of different methods on Vicuna-7B and 13B models. Efficiency improvements of MoA are ablated with four factors. All sparse attention methods use 50% density. Decode throughput (tokens per second) evaluated at the maximum batch capacity of an A100-80GB GPU.
4k 8k 16k
Model Framework Attention Batch Throughput Batch Throughput Batch Throughput
7B vLLM PagedAttention 30 628.8 15 323.0 8 145.5
FlexGen H2O 20 754.9 6 296.3 1 51.7
HuggingFace InfLLM 15 62.0 10 37.5 6 19.2
HuggingFace StreamingLLM 50 945.1 25 467.3 12 232.0
HuggingFace FlashAttention2 30 134.6 15 66.9 8 32.9
+Static KV-Cache 30 496.1 15 219.5 8 91.6
+Reduced Attention 30 722.5 15 369.9 8 178.3
+Increased Batch 50 897.7 25 436.7 12 206.4
+Kernel (=MoA) 50 1099.0 25 535.7 12 257.3
13B vLLM PagedAttention 16 314.8 8 160.5 4 71.1
FlexGen H2O 12 330.2 4 138.2 1 37.4
HuggingFace InfLLM 8 30.3 5 17.63 3 11.3
HuggingFace StreamingLLM 28 478.4 14 241.2 7 116.5
HuggingFace FlashAttention2 16 81.3 8 40.8 4 19.8
+Static KV-Cache 16 264.6 8 111.3 4 62.2
+Reduced Attention 16 329.6 8 156.4 4 87.3
+Increased Batch 28 471.5 14 222.6 7 108.3
+Kernel (=MoA) 28 550.9 14 267.6 7 132.3
MoA shows high runtime efficiency with a manageable one-time compression overhead.
Runtime Efficiency. Table 5 compares the runtime efficiency of MoA over various attention methods and LLM frameworks, with the ablation of efficiency improvements brought by each design factor of MoA. At 50% density, MoA boosts the decode throughput by 6.6$\times$ to 8.2$\times$ compared to FlashAttention2. It outperforms H2O and InfLLM with 1.2×1.2$\times$1.2 × to 4.0×4.0$\times$4.0 × decode throughput improvements. Even compared to the highly system-level optimized vLLM framework (Kwon et al., 2023), MoA still achieves a 1.7×1.7$\times$1.7 × to 1.9×1.9$\times$1.9 × throughput increase. MoA also reduces total GPU memory by 1.2×1.2$\times$1.2 × to 1.4×1.4$\times$1.4 ×, are detailed in Appendix B.2. This throughput gain results from four main factors: static-sized KV-Cache during generation(3.0$\approx$ 3.0$\times$ 3.0 ×); reduced attention computations due to sparsity (1.5$\approx$ 1.5$\times$ 1.5 ×); increased batch sizes enabled by smaller KV-Cache memory (1.4$\approx$ 1.4$\times$ 1.4 ×); and our CUDA-implemented GPU kernel for MoA heterogeneous attention (1.2$\approx$ 1.2$\times$ 1.2 ×).
Compression Pipeline Efficiency. MoA completes the automatic compression pipeline for the Vicuna-7B and 13B models within two hours. For the larger Llama3-70B model, the process requires 8.5 hours of real-time and 34.7 hours of GPU time. Details for each compression step overhead are in Appendix B.2.2.
@@ 6.5 Rules Discovered by MoA
We investigate MoA’s elastic rules for each head. As shown in Figure 10, masks in the initial and middle layers exhibit high density, aligning with the conclusions from previous research on LLM’s intrinsic dimensions (Valeriani et al., 2023) and layer sensitivities (Yuan et al., 2023). Conversely, in the final layers, most heads require low density, while few need high density. Figure 11 shows that layers with lower average mask density typically display a wider range of densities among heads, confirming the need for heterogeneity within the same layer. Further details and quantified insights are presented in Appendix C.
@@ 7 Conclusion and Future Work
MoA automates the selection of heterogeneous elastic masks for each attention head and input length, significantly extending the effective context length of LLMs by 3.9×3.9$\times$3.9 ×. It enhances retrieval accuracy by 1.5×1.5$\times$1.5 × to 7.1×7.1$\times$7.1 × over uniform sparse attention method and increases throughput to over 6×6$\times$6 × at 50% average density, maintaining performance on par with dense models in rigorous benchmarks.
Limitations and Future Work. Under an extremely low-density budget, MoA fails to maintain good performance. Designing a dynamic MoA method has the potential to address this issue, which we leave for future work. Using non-linear elastic rules with bounded attention spans is also worth exploring. Additionally, MoA’s profiling method can be adapted to evaluate the influence of weights and other activations, facilitating other compression methods such as weight and activation quantization (Lin et al., 2023; Xiao et al., 2024b; Li et al., 2024b).
@@ A.1 Main Setup
Baselines. In the setup for our experiment, we adhere to specific configurations outlined in the respective papers. In the case of In StreamingLLM (Xiao et al., 2024c), the initial four tokens remain unmasked, serving as the attention sink, except for the 70b model in Table 4 and the super long setting in Figure 5, where we use 64 tokens as the attention sink. For InfLLM (Xiao et al., 2024a), we adhere to the original configuration by maintaining the same local window size and selected memory size, using 128 initial tokens as specified in their setup. For H2O (Zhang et al., 2023), we ensure the same number of heavy hitter tokens and recent tokens. Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction. StreamingLLM, InfLLM and MoA use sparse prefill.
@@
Models and Benchmarks. Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k (Chiang et al., 2023) can only take in 16k context length, we use the 16k split of LV-Eval benchmark (Yuan et al., 2024), truncating the input to 15500 for model input in Table 4. For the LongBench benchmark (Bai et al., 2023), we use the LongBench-E split, which features a balanced number of data items at every length level. The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens. We test each split using the input length truncation thresholds of 3,500, 7,500, and 15,500 tokens, respectively.
@@
Perplexity Evaluation. We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016), representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM. Following LongBench, the data items are organized as question-answer pairs. The questions and answers are written by humans and come with the dataset. The perplexity is calculated solely on the answer part of the data, demonstrating the model’s coherence in responding to user requests.
@@
Validation Dataset. The validation dataset is used to select the optimal compression plan among the Pareto front solutions during the optimization step. The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets. $50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016). The additional $50$ data items from the LongEval (Li et al., 2023) dataset are also added to validate the retrieval ability. For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
@@
MoA Settings. MoA uses the block sparse attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block. The first block of tokens is not masked as the attention sink. For the profile stage, we use the MultiNews (Fabbri et al., 2019) calibration dataset with model response as supervision, as described in Section 5. We use $50\times 3$ data items at 2k, 4k, 8k lengths. The data items are padded to their corresponding length level in order to ensure a unified shape of attention influence tensors for each length level. We adopt block granularity during the profiling stage, calculating the average attention influence within each block to represent the block’s overall influence. The optimization is done with the multi-objective optimization at the same set of lengths. We limit the number of distinct rules to at most two per model layer to ensure inference-time efficiency. Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
@@ A.2 Efficiency Experiment Setup
For MoA, The implementation is based on Huggingface Transformers. During the prefill stage, we use the sparse CUDA kernel designed by us with block size 64. During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules. Thanks to our fixed sliding-window span during the decode stage, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache. Our custom decoding CUDA kernel then handles KV-Cache with varying lengths across different attention heads during the decoding process.
For H2O, we use its official efficient implementation, which is based on Flexgen (Sheng et al., 2023). Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated. It makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention. Therefore, H2O is easy to get OOM (Out-Of-Memory) with large prefill length and increased batch size.
@@
Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible. However, the official efficient implementation of H2O based on Flexgen only supports OPT (Zhang et al., 2022). Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 9 for comparison.